OpenStudy (anonymous):

Find the cot^-1(-1/2) if tan^-1(-2) is about -63 degrees. explain.

OpenStudy (dumbcow):

could you clarify the expression please...is that inverse cotangent

OpenStudy (anonymous):

yes that is sorry I made a typo

OpenStudy (dumbcow):

what do you mean by "of" ?

OpenStudy (anonymous):

yes that is correct

OpenStudy (dumbcow):

oh ok , if

OpenStudy (anonymous):

i don't does it mean times?

OpenStudy (dumbcow):

you should write your question as: Given tan^-1 (-2) = 63 degrees Find cot^-1 (-1/2) \[\tan(\theta) = \frac{y}{x} \rightarrow \cot (\theta) = \frac{x}{y}\] thus \[\tan^{-1} (\frac{y}{x}) = \cot^{-1} (\frac{x}{y}) \] -1/2 is reciprocal of -2, therefore it is also equal to about -63 degrees

OpenStudy (anonymous):

goody eye!

OpenStudy (anonymous):

Oh okay that makes sense thank you! my teacher was also talking about restrictions too which I don't understand where they come in to this problem

OpenStudy (dumbcow):

well since you are using tan, theta can't be pi/2

OpenStudy (anonymous):

Okay so I would say there is a restiction at pi/2?